# Rethinking Approximate Gaussian Inference in Classification

### v1

@@ C.2 Analysis of the Probit Approximation for Sigmoid Models
    it tends to 0 as $\delta$ tends to 0.
In other words, given knowledge of the worst case error in the approximation Equation 49 on the compact set $\mathcal{K}$, we can bound the KL divergence in terms of that error independently of the number of classes. Due to the simplicity of our assumptions, the bound remains quite raw and could be strengthened with further distributional assumptions on the means and variances.
Proof.
$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})$ $=\sum_{c=1}^{C}p_{c}(\log p_{c}-\log\hat{p}_{c})$ (52)
$=\sum_{c=1}^{C}\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\left(\log\left(\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)-\log\left(\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}}\right)\right)$
$\leq\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(\frac{q_{c}-\delta}{q_{c}}\right)+\log\left(\frac{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}+C\delta}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)\right)$
$=\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(1-\frac{\delta}{q_{c}}\right)+\log\left(1+\frac{C\delta}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)\right)$
$\leq\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(1-\frac{\delta}{u}\right)+\log\left(1+\frac{C\delta}{Cu}\right)\right)$
$=\log\left(\frac{u+\delta}{u-\delta}\right).$
C.3 Analysis of the Mean Field Approximation for Softmax Models
Recall the mean field approximation to the Gaussian softmax integral (Section A.4)
$\bm{p}:=\int_{\mathbb{R}^{C}}\operatorname{\mathbf{softmax}}(\bm{x})\;\mathcal{N}(\bm{\mu},\bm{\Sigma})(d\bm{x})\approx\operatorname{\mathbf{softmax}}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)=:\bm{\hat{p}}.$ (53)
This approximation is obtained through three subsequent approximations, that is, (1), (2), and (3) in Equation 36. Approximation (2) is not actually necessary, as it can be replaced by an exact computation, as shown in (Lu et al., 2021). However, this empirically does not appear to improve the quality of the overall approximation. We focus our attention on the main approximation (1), namely
$\mathbb{E}\left[\left(2-C+\sum_{c^{\prime}\neq c}\rho(Y_{c}-Y_{c^{\prime}})^{-1}\right)^{-1}\right]\stackrel{{\scriptstyle(1)}}{{\approx}}\left(2-C+\sum_{c^{\prime}\neq c}\mathbb{E}[\rho(Y_{c}-Y_{c^{\prime}})]^{-1}\right)^{-1}.$ (54)
This is exact if $\rho(Y_{c}-Y_{c^{\prime}})$ is a constant random variable for all $Y_{c}$ and $Y_{c^{\prime}}$. For $C>2$, this would imply that $Y_{c}$ is distributed according to a Dirac delta. However, this does not allow for any second-order distribution over the logits. We thus see that we cannot train for exactness in the mean field approximation as we do for Equation 13. This is a fundamental limitation of this mean field approximation.
@@ Appendix D Moment Matching Beta distributions
by moment matching. Specifically, the parameters $\bm{\alpha},\bm{\beta}\in\mathbb{R}_{>0}^{C}$ that match the moments of $\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}$ for some $\mathrm{f}\in\mathcal{G}(\mathbb{R}^{C})$ must satisfy
@@
where all vector operations are element-wise. Multiplying out the denominators on the right-hand side of the equations of Equation 56, we obtain a system of two linear equations with two unknowns (for each $c$), which can be solved uniquely, yielding
@@ D.1 Information Geometric Interpretation of the Pushforward through NormCDF
Figure 10: Illustration of the statistical manifolds of the pushforward Gaussian distributions $\mathcal{G}(\mathbb{R})$ through the normCDF and the sigmoid respectively compared to the statistical manifold of Beta distributions $\mathcal{B}((0,1))$. NormCDF, unlike the sigmoid, makes the manifolds intersect at the point $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$.
Here, we argue that the normCDF activation is an ideal choice for approximating Gaussian pushforwards with Beta distributions by interpreting Figure 10. This extends the work from (MacKay, 1998), as it shows how one can make sense of the ‘right’ basis for performing Laplace approximations in the classification setting.
@@
The key observation is that $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$ are isometric to $\mathcal{G}(\mathbb{R})$. This is because $\Phi$ and $\rho$ are diffeomorphisms, so in particular sufficient statistics, and the Fisher information metric is invariant under sufficient statistics (Ay et al., 2017, Section 5.1.3). Visually, this means that $\varphi_{*}\mathcal{G}(\mathbb{R})$ has the same shape irrespectively of the diffeomorphism activation function $\varphi$. One can thus observe that, given that $\mathcal{B}((0,1))$ is not isometric to $\mathcal{G}(\mathbb{R})$, there exists no activation $\varphi$ such that $\varphi_{*}\mathcal{G}(\mathbb{R})=\mathcal{B}((0,1))$. To design an activation $\varphi$ that maps Gaussians to Betas, the best one can hope to do is to map one specific Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ to a specific Beta distribution $B(\alpha,\beta)$. This is done with the map $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}$ where $\Phi_{\mu,\sigma^{2}}$ and $F_{\alpha,\beta}$ are the cumulative distribution functions of $\mathcal{N}(\mu,\sigma^{2})$ and $B(\alpha,\beta)$ respectively. Taking $\mu=0$, $\sigma^{2}=1$, $\alpha=\beta=1$ we get $\Phi_{0,1}=\Phi$ and $F_{1,1}=\operatorname{id}_{(0,1)}$, yielding $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}=\Phi$.
Now $\mathcal{G}(\mathbb{R})$, and hence $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$, is isometric to the hyperbolic plane Ay et al. (2017, Example 3.1). We can embed this isometrically into Minkowski space with the classical hyperboloid model of the hyperbolic plane (Reynolds, 1993),
@@
For $\mathcal{B}((0,1))$, we use the isometric embedding from Le Brigant et al. (2021, Proposition 2):
@@
Finally, we choose our embedding Equation 58 for $\Phi_{*}\mathcal{G}(\mathbb{R})$ such that it intersects the embedding Equation 59 at a point, to highlight that the statistical manifolds $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ intersect at a point in the infinite-dimensional ambient space $\mathcal{M}((0,1))$, while $\rho_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ do not.
Moreover, since $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$, our moment matching approximation Equation 56 is exact when $\mathrm{f}(x)$ is the standard Normal distribution.
Appendix E Comparison with the Energy Model View of Classifiers
Grathwohl et al. (2019) propose to view classifiers of the form Equation 2 as energy-based models by
$p(c\mid x)$ $=\frac{\exp(f_{c}(x))}{\sum_{c^{\prime}=1}^{C}\exp(f_{c^{\prime}}(x))},$ (60)
$p(x,c)$ $=\frac{\exp(f_{c}(x))}{Z},$
$p(x)$ $=\frac{\sum_{c=1}^{C}\exp(f_{c}(x))}{Z}$
where $Z=\int_{\mathcal{X}}\sum_{c=1}^{C}\exp(f_{c}(x))\;dx$. Note the consistency
$p(c\mid x)=\frac{p(x,c)}{p(x)}.$ (61)
Observe that the choice of $p(x,c)$ and $p(x)$ in Equation 60 is not canonical: we could have multiplied $p(x,c)$ and $p(x)$ by $\frac{h(x)}{Z^{\prime}}$ for any non-vanishing function $h$ and appropriate normalisation constant $Z^{\prime}$, and we would still obtain the consistency Equation 61. Grathwohl et al. (2019) propose to use the degree of freedom in the logits to train for Equation 60, turning discriminative models into generative models.
Note that $\exp$ in Equation 60 can be replaced by any positive activation $\varphi$. In our framework, we train for $\sum_{c=1}^{C}\varphi(f_{c}(x))=s(x)$. More specifically, in the present work, we train for $\sum_{c=1}^{C}\varphi(f_{c}(x))=1$ (Section 4), making use of the degree of freedom in the logits to obtain exact predictives in Equation 13. In the formalism Equation 60, this means we are setting $x$ to be uniformly distributed. An interesting avenue of future research would be to combine our framework with that of Grathwohl et al. (2019), by e.g. training for $\sum_{c=1}^{C}\varphi(f_{c}(x))\propto p(x)$, and extending the framework of Grathwohl et al. (2019) to second-order distributions.
Appendix F List of Predictive and Dirichlet Parameters Formulas
F.1 Predictive Formulas
@@ F.1.1 Softmax
The Mean Field predictive (Lu et al., 2021) uses the following approximation for the true predictive:
@@
The Laplace Bridge predictive (Hobbhahn et al., 2022) approximates the true predictive as follows:
@@
F.1.2 Exp
Our analytic predictive for the exp activation function (Section 2.1) is given by
@@ F.1.2 Exp
F.1.3 NormCDF
The analytic predictive for the normCDF activation function (Section 2.2) is computed as
@@ F.1.3 NormCDF
F.1.4 Sigmoid
For the sigmoid activation function (Section 2.2), the analytic predictive can be computed as
@@ F.1.4 Sigmoid
F.2 Dirichlet Parameters Formulas
@@
F.2.1 Softmax
The Laplace Bridge method (Hobbhahn et al., 2022) uses the following Dirichlet parameters:
@@ F.2.1 Softmax
F.2.2 Exp
Exp uses the analytic parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
@@ F.2.2 Exp
F.2.3 NormCDF
NormCDF uses the analytic parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
@@ F.2.3 NormCDF
F.2.4 Sigmoid
Sigmoid uses the analytic parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
@@ F.2.4 Sigmoid
$\bm{a}=\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right),\;\bm{b}=\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)-\frac{1}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)\left(1-\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)\right).$ (76)
Appendix G List of Uncertainty Estimators
@@ G.1 Predictive
G.2 Monte Carlo
Given $S$ Monte Carlo samples $\bm{\hat{p}}^{(1)},\dots,\bm{\hat{p}}^{(S)}$ with mean $\bm{\hat{p}}$, one can calculate a predictive as their average and derive the estimators in Section G.1. However, Monte Carlo samples allow one to calculate two additional estimators detailed below.
@@ G.3 Dirichlet
Appendix H Experimental Setup
@@
To answer the first question, we evaluate our analytic predictives (Sigmoid, NormCDF, Exp) and moment-matched Dirichlet distributions against softmax models equipped with approximate inference tools (Laplace Bridge (Hobbhahn et al., 2022), Mean Field (Lu et al., 2021), Monte Carlo sampling). We consider Heteroscedastic Classifiers (HET) (Collier et al., 2021), Spectral-Normalized Gaussian Processes (SNGP) (Liu et al., 2020), and last-layer Laplace approximation methods (Daxberger et al., 2021) as backbones (see Appendix J for details).
The resulting 18 (method, activation, predictive) triplets are evaluated on ImageNet-1k (Deng et al., 2009) and CIFAR-10 (Krizhevsky & Hinton, 2009) on five metrics aligning with practical needs from uncertainty estimates (Mucsányi et al., 2023):
@@
See Appendix I for details.
For ImageNet, we treat ImageNet-C (Hendrycks & Dietterich, 2019) samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
@@
For ImageNet evaluation, we use a ResNet-50 backbone pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following Mucsányi et al. (2024). We train with the LAMB optimizer (You et al., 2019) using a batch size of $128$ and gradient accumulation across $16$ batches, resulting in an effective batch size of $2048$, following Tran et al. (2022). We further use a cosine learning rate schedule with a single warmup epoch using a warmup learning rate of $0.0001$. The learning rate is treated as a hyperparameter and selected from the interval $[0.0005,0.05]$ based on the validation performance. The weight decay is selected from the set $\{0.01,0.02\}$. During training, we keep track of the best-performing checkpoint on the validation set and load it before testing. We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme (Shahriari et al., 2015) in Weights & Biases (Biewald, 2020) based on the negative log-likelihood.
On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs. The only exceptions are the SNGP models that are trained for 125 epochs (Liu et al., 2020). We train with Momentum SGD using a batch size of $128$ and no gradient accumulation. Similarly to ImageNet, we use a cosine learning rate schedule but with five warmup epochs and warmup learning rate $1\mathrm{e}{-5}$. The learning rate is also treated as a hyperparameter on CIFAR-10. We use the interval $[0.05,1]$ for Sigmoid and NormCDF, and $[0.01,0.15]$ for Softmax and Exp. The optimal learning rates for Sigmoid and NormCDF are generally larger, as the class-wise binary cross-entropies are averaged instead of summed. The weight decay is selected from the interval $[1\mathrm{e}{-6},1\mathrm{e}{-4}]$. Similarly to ImageNet, we use the best-performing checkpoint in the tests and use a ten-step Bayesian Optimization scheme to select performant hyperparameters.
Appendix I Benchmark Metrics

### v2

@@ C.1 Informal Analysis of Monte Carlo Approximations
    $M(\mathcal{K})\to 0$ as $\delta\to 0$.
In other words, given knowledge of the worst case error in the approximation Eq. 15 on the compact set $\mathcal{K}$, we can bound the KL divergence in terms of that error independently of the number of classes. Due to the simplicity of our assumptions, the bound remains quite raw and could be strengthened with further distributional assumptions on the means and variances.
Finally, to obtain a meaningful bound in Theorem C.2, we needed to assume $\Delta(\mathcal{K}):=\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$. In Fig. C.2 we see that this can be assumed to hold on the compact sets on which logit means and variances tend to live (see also Fig. C.1).
Figure C.2: Plot of the ‘true’ predictive $q$ (approximated with a $50000$ sample MC approximation) versus the probit approximation $\hat{q}$, as well as $\frac{q-\hat{q}}{q}$. We see that $\frac{q-\hat{q}}{q}$ tends to be negative, making the assumption $\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$ from Theorem C.2 easily fulfilled.
@@ Appendix D Moment Matching Beta Distributions
by moment matching. Specifically, the parameters $\bm{\alpha},\bm{\beta}\in(0,\infty)^{C}$ that match the moments of $\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}$ for some $\mathrm{f}\in\mathcal{G}(\mathbb{R}^{C})$ must satisfy
@@
where all vector operations are element-wise. Multiplying out the denominators on the right-hand side of the equations of Eq. 60, we obtain a system of two linear equations with two unknowns (for each $c$), which can be solved uniquely, yielding
@@ D.1 Information Geometric Interpretation of the Pushforward through NormCDF
Figure D.1: Illustration of the statistical manifolds of the pushforward Gaussian distributions $\mathcal{G}(\mathbb{R})$ through the normCDF and the sigmoid respectively compared to the statistical manifold of Beta distributions $\mathcal{B}((0,1))$. NormCDF, unlike the sigmoid, makes the manifolds intersect at the point $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$.
Here, we argue that the normCDF activation is an ideal choice for approximating Gaussian pushforwards with Beta distributions by interpreting Fig. D.1. This extends the work from [22], as it shows how one can make sense of the ‘right’ basis for performing Laplace approximations in the classification setting.
@@
The key observation is that $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$ are isometric to $\mathcal{G}(\mathbb{R})$. This is because $\Phi$ and $\rho$ are diffeomorphisms, so in particular sufficient statistics, and the Fisher information metric is invariant under sufficient statistics [2, Section 5.1.4]. Visually, this means that $\varphi_{*}\mathcal{G}(\mathbb{R})$ has the same shape irrespectively of the diffeomorphism activation function $\varphi$. One can thus observe that, given that $\mathcal{B}((0,1))$ is not isometric to $\mathcal{G}(\mathbb{R})$, there exists no activation $\varphi$ such that $\varphi_{*}\mathcal{G}(\mathbb{R})=\mathcal{B}((0,1))$. To design an activation $\varphi$ that maps Gaussians to Betas, the best one can hope to do is to map one specific Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ to a specific Beta distribution $B(\alpha,\beta)$. This is done with the map $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}$ where $\Phi_{\mu,\sigma^{2}}$ and $F_{\alpha,\beta}$ are the cumulative distribution functions of $\mathcal{N}(\mu,\sigma^{2})$ and $B(\alpha,\beta)$ respectively. Taking $\mu=0$, $\sigma^{2}=1$, $\alpha=\beta=1$ we get $\Phi_{0,1}=\Phi$ and $F_{1,1}=\operatorname{id}_{(0,1)}$, yielding $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}=\Phi$.
Now $\mathcal{G}(\mathbb{R})$, and hence $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$, is isometric to the hyperbolic plane Ay et al. [2, Example 3.1]. We can embed this isometrically into Minkowski space with the classical hyperboloid model of the hyperbolic plane [31],
@@
For $\mathcal{B}((0,1))$, we use the isometric embedding from Le Brigant et al. [18, Proposition 2]:
@@
Finally, we choose our embedding Eq. 62 for $\Phi_{*}\mathcal{G}(\mathbb{R})$ such that it intersects the embedding Eq. 63 at a point, to highlight that the statistical manifolds $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ intersect at a point in the infinite-dimensional ambient space $\mathcal{M}((0,1))$, while $\rho_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ do not.
Moreover, since $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$, our moment matching approximation Eq. 60 is exact when $\mathrm{f}(x)$ is the standard Normal distribution.
Appendix E List of Predictive and Dirichlet Parameters Formulas
E.1 Predictive Formulas
@@ E.1.1 Softmax
The Mean Field predictive [21] uses the following approximation for the true predictive:
@@
The Laplace Bridge predictive [14] approximates the true predictive as follows:
@@
E.1.2 Exp
Our closed-form predictive for the exp activation function (Section 2.1) is given by
@@ E.1.2 Exp
E.1.3 NormCDF
The closed-form predictive for the normCDF activation function (Section 2.2) is computed as
@@ E.1.3 NormCDF
E.1.4 Sigmoid
For the sigmoid activation function (Section 2.2), the closed-form predictive can be computed as
@@ E.1.4 Sigmoid
E.2 Dirichlet Parameters Formulas
@@
E.2.1 Softmax
The Laplace Bridge method [14] uses the following Dirichlet parameters:
@@ E.2.1 Softmax
E.2.2 Exp
Exp uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
@@ E.2.2 Exp
E.2.3 NormCDF
NormCDF uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
@@ E.2.3 NormCDF
E.2.4 Sigmoid
Sigmoid uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
@@ E.2.4 Sigmoid
$\bm{a}$ $=\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right),$ (78)
$\bm{b}$ $=\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)-\frac{1}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)\left(1-\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)\right).$
Appendix F List of Uncertainty Estimators
@@ F.1 Predictive
F.2 Monte Carlo
Given $S$ Monte Carlo samples $\bm{\hat{p}}^{(1)},\dots,\bm{\hat{p}}^{(S)}$ with mean $\bm{\hat{p}}$, one can calculate a predictive as their average and derive the estimators in Section F.1. However, Monte Carlo samples allow one to calculate two additional estimators detailed below.
@@ F.3 Dirichlet
Appendix G Experimental Setup
@@
To answer the first question, we evaluate our closed-form predictives (Sigmoid, NormCDF, Exp) and moment-matched Dirichlet distributions against softmax models equipped with approximate inference tools (Laplace Bridge [14], Mean Field [21], Monte Carlo sampling). We consider Heteroscedastic Classifiers (HET) [5], Spectral-Normalized Gaussian Processes (SNGP) [20], and last-layer Laplace approximation methods [8] as backbones (see Appendix I for details).
The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [9] and CIFAR-10 [17] on five metrics aligning with practical needs from uncertainty estimates [26]:
@@
See Appendix H for details.
For ImageNet, we treat ImageNet-C [13] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
@@
For ImageNet evaluation, we use a ResNet-50 backbone pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following Mucsányi et al. [27]. We train with the LAMB optimiser [40] using a batch size of $128$ and gradient accumulation across $16$ batches, resulting in an effective batch size of $2048$, following Tran et al. [36]. We further use a cosine learning rate schedule with a single warmup epoch using a warmup learning rate of $0.0001$. The learning rate is treated as a hyperparameter and selected from the interval $[0.0005,0.05]$ based on the validation performance. The weight decay is selected from the set $\{0.01,0.02\}$. During training, we keep track of the best-performing checkpoint on the validation set and load it before testing. We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [33] in Weights & Biases [3] based on the negative log-likelihood.
On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs. The only exceptions are the SNGP models that are trained for 125 epochs [20]. We train with Momentum SGD using a batch size of $128$ and no gradient accumulation. Similarly to ImageNet, we use a cosine learning rate schedule but with five warmup epochs and warmup learning rate $1\mathrm{e}{-5}$. The learning rate is also treated as a hyperparameter on CIFAR-10. We use the interval $[0.05,1]$ for Sigmoid and NormCDF, and $[0.01,0.15]$ for Softmax and Exp. The optimal learning rates for Sigmoid and NormCDF are generally larger, as the class-wise binary cross-entropies are averaged instead of summed. The weight decay is selected from the interval $[1\mathrm{e}{-6},1\mathrm{e}{-4}]$. Similarly to ImageNet, we use the best-performing checkpoint in the tests and use a ten-step Bayesian Optimization scheme to select performant hyperparameters.
The hyperparameter optimization, training, and evaluation of the methods used in this paper took $0.8$ GPU years on NVIDIA RTX 2080Ti GPUs in a university compute cluster. The individual runs required no more than $50$ GB of RAM and $3$ days of runtime.
Appendix H Benchmark Metrics